# Prompt-SAW: Leveraging Relation-Aware Graphs for Textual Prompt Compression

### v1

@@ Abstract
Large language models (LLMs) have shown exceptional abilities for multiple different natural language processing tasks. While prompting is a crucial tool for LLM inference, we observe that there is a significant cost associated with exceedingly lengthy prompts. Existing attempts to compress lengthy prompts lead to sub-standard results in terms of readability/interpretability of the compressed prompt, with a detrimental impact on prompt utility. To address this, we propose Prompt-SAW: Prompt compresSion via Relation AWare graphs, an effective strategy for prompt compression over task-agnostic and task-aware prompts. Prompt-SAW uses the prompt’s textual information to build a graph, later extracts key information elements in the graph to come up with the compressed prompt. We also propose GSM8k-aug, i.e., an extended version of the existing GSM8k benchmark for task-agnostic prompts in order to provide a comprehensive evaluation platform. Experimental evaluation using benchmark datasets shows that prompts compressed by Prompt-SAW are not only better in terms of readability, but they also outperform the best-performing baseline models by up to 14.3% and 13.7% respectively for task-aware and task-agnostic settings while compressing the original prompt text by 33.0% and 56.7%.
@@ 1 Introduction
Large Language Models (LLMs) have attracted considerable attention for their superior performance across a wide range of applications. For this, instructions (aka. prompts) play a crucial role in extending the capabilities of LLMs for multiple different tasks. The prompts provide the provision to guide the model to elucidate desired model behavior without perturbing the model parameters. This is also highlighted in recent studies that show well-designed prompts and the integration of external knowledge are significant to enhance the effectiveness of LLMs’ (Sahoo et al., 2024). Different LLMs-related techniques directly benefiting from prompts include but are not limited to: In-Context Learning (Dong et al., 2022), Chain-of-Thought (Wei et al., 2022), Retrieval Augmented Generation (Lewis et al., 2020), and Agents (Park et al., 2023) etc. Generally, prompts may be sub-divided into two types: task-aware and task-agnostic prompts, a quick overview is given in Appendix A.2 and Appendix A.3 respectively.
At the same time, the abilities of LLMs are significantly compromised/constrained by increasingly lengthy prompts, even comprising thousands of tokens. Lengthy prompts not only obscure requisite information but also increase computational costs and incur inference latency. To tackle this challenge, prompt compression techniques, e.g., (Li, 2023), have garnered significant interest. These approaches are based on the fact that natural language is inherently redundant (Shannon, 1951). Thus, it is possible to substantially compress the length of original textual prompts by preserving requisite information in small segments.
@@
While existing approaches could enhance the ability to deal with lengthy prompts for LLMs, they lack grammatical coherence, i.e., existing approaches neglect the syntactic and semantic structure of the compressed prompt. This is because contemporary prompt compression methods primarily focus on quantifying token-level information, neglecting the overall grammatical structure of the compressed prompt. Such ignorance not only increases the risk of semantic loss within the compressed prompt but also hampers its readability for human readers. An example in this regard is shown in Figure 1, where the original prompt text: ”Two women have won the prize: Curie and Maria Goeppert-Mayer” is compressed to: ”won twoes forg01 theate women prize:ertMayer” by LongLLMlingua Jiang et al. (2023b).
Figure 1: An exmple of prompt compression example of previous token-level based method LongLLMlingua (Jiang et al., 2023b)
@@
We conducted extensive experimental analysis of Prompt-SAW under both task-agnostic and task-aware settings against existing best-performing models as baselines. For evaluation, we used: (i)  GSM8k-aug, i.e., an extended experimental setting proposed by us for open-source benchmark GSM8K (Cobbe et al., 2021), and (ii) open-source benchmark NaturalQuestions (Liu et al., 2023). Experimental results show that Prompt-SAW significantly outperforms other baseline models. We summarize the key contributions of this work as follows:
@@
  • As current benchmarks for task-agnostic prompts lack comprehensive evaluation, we propose GSM8k-aug, an extended version of existing GSM8k benchmark for an intensive evaluation of Prompt-SAW.
  • We demonstrated the effectiveness of Prompt-SAW by comprehensive experiments showing Prompt-SAW attained state-of-the-art performance outperforming baseline models by up to 14.3% and 13.7% respectively for task-aware and task-agnostic settings while compressing the original prompt text by 33.0% and 56.7%.
@@ 2 Related Work
Soft prompts were introduced by Lester et al. (2021). A soft prompt integrates additional trainable parameters at the model’s input stage. (Wingate et al., 2022) emphasized that soft prompt compression effectively retains crucial abstract information with a reduced parameter count. Xu et al. (2023b) emphasized that carefully crafted prompts are helpful in augmenting the end-performance of compressed LLMs, also the compressed LLMs are helpful in the prompt learning phase.
Compared to soft prompt compression, discrete prompt compression seeks to optimize the effectiveness of prompts via token-level search strategies. Jung & Kim (2023) employed policy networks to eliminate unnecessary tokens for prompt compression. Li (2023) utilized self-information metrics to identify and remove superfluous information in prompts. Capitalizing on these advancements, Jiang et al. (2023a) and Jiang et al. (2023b) formulated algorithms for dynamically adjusting compression rates across different prompt sections, giving precedence to tokens with higher perplexity.
@@ 3 Preliminaries
We use $P$ and $C$ to represent the original and compressed prompt respectively. Likewise, we use $N$ and $\widetilde{N}$ to represent the length of the original and compressed prompt. We use $\eta=\widetilde{N}/N$ to represent the compression rate and $1/\eta$ as the compression ratio. $\eta^{*}$ is used to represent the target compression rate. The graph is represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $e_{i}$, $r_{i}$ and $e_{i}^{{}^{\prime}}$ represent the subject entity, relation and object entity in the graph respectively; =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ and =$\mathcal{R}=\{r_{1},r_{2},\cdots,r_{n}\}$ denote the set of entities and relations in $\mathcal{G}$. $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is used to represent small-scale information elements in $\mathcal{G}$, this is equivalent to graph triplet. $M$ represents auxiliary models used for graph construction. $E$ is used to represent the encoder network. $\delta$ represents the similarity threshold used for sub-graph construction.
@@ 4.1 Motivation of Prompt-SAW
Formally, given a prompt text $P$, we claim it encompasses a set of entities $(\mathcal{E})$, i.e., names of persons, locations, organizations, miscellaneous elements, etc., (Ali et al., 2020; 2021). These entities serve as the key elements of the prompt structure. In addition to the entities, we can also infer some relations $\mathcal{R}$ in $P$ that may be used to describe the connections between the entities. Prompt-SAW re-organizes these key elements of the prompt (i.e., entities and their relations) in a graph structure, represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$. We use $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ to represent the $i$-th information element of $\mathcal{G}$, i.e., a fact stating that $e_{i}$ has $r_{i}$-$th$ relation with $e_{i}^{{}^{\prime}}$.
@@ 4.2 Workflow of Prompt-SAW
The workflow of Prompt-SAW consists of two parts. First, it uses the information in prompt $P$ to construct a graph $\mathcal{G}$. Then, based on the specific scenario:
@@
We will provide details in the following subsections.
4.2.1 Graph Construction.
For graph construction from the text data, we use an in-context learning prompt that prompts a language model to construct the graph from the original prompt text as follows.
$\mathcal{G}={M}(P_{\text{template}}(P)),$ (1)
where $(P)$ is the prompt text and $P_{\text{template}}$ is the prompt template (explained in Appendix B.1) used to guide the model (M) to extract the graph structure $(\mathcal{G})$. For graph construction, we typically prefer a small-scale model for $M$, as a lesser number of parameters can avoid a higher computational cost incurred by large models, which eventually contributes to the overall compression time overhead. It is notable that in addition to the above-mentioned method, we also employed traditional approaches for graph construction, e.g., OpenIE (Pal & Mausam, 2016), however, these approaches yield relatively lower performance.
4.2.2 Task-aware Prompts
$Emb_{p^{\mathrm{que}}}=E(p^{\mathrm{que}})$ (2)
$Index_{\text{ranked}}=\text{Rank}(Sim_{\mathcal{G}})$ (4)
$C=e_{1}\oplus r_{1}\oplus e_{1}^{\prime}\oplus\text{[SEP]}\oplus\cdots e_{i}\oplus r_{i}\oplus e_{i}^{\prime}\oplus\text{[SEP]}\oplus\cdots e_{n}\oplus r_{n}\oplus e_{n}^{\prime}$ (5)
where $\oplus$ is the concatenation operator used to combine the entities and relations within the information elements $(g_{i})$ in the extracted subgraph $\mathcal{G}_{\text{subset}}$, and [SEP] is the delimiter used to separate different information elements in $\mathcal{G}_{\text{subset}}$.
4.2.3 Task-agnostic Prompts
Compress Function. The workflow of the $Compress()$ is shown in Appendix C and explained as follows: (i) start with an empty graph $(\mathcal{G}^{\prime})$ (line-1); (ii) iterate the information elements in the graph $(g_{i}\in\mathcal{G})$ to compute the similarity score of $g_{i}$ with the elements in $\mathcal{G}^{\prime}$ to look for maximal similarity, i.e., $sim_{max}$ (lines 2-3); (iii) compare $sim_{max}$ against the compression threshold $\delta$ to insert $g_{i}$ in $(\mathcal{G}^{\prime})$ (lines 4-5); (iv) finally, return $(\mathcal{G}^{\prime})$ as the final subset of the graph. The end-goal of Algorithm 2 is to select highly dis-similar information elements (Ali et al., 2019) by neglecting cases with $sim_{max}>\delta$. For such cases, we assume that the corresponding information element, i.e., $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is redundant because there is already an element in $\mathcal{G}^{\prime}$ that is very similar to $g_{i}$.
5 Experiments
5.1 Experiment Settings
Datasets. To comprehensively evaluate the effectiveness of compressed prompts, we evaluate their performance on task-agnostic and task-aware datasets. For task-agnostic dataset, we propose GSM8k-aug, i.e., an extended variant of the original GSM8K (Cobbe et al., 2021) devised by us to report the model performance under $i$-shot settings with $i\in\{1,2,4\}$ (see Appendix D.1 for details). Note, GSM8k-aug has a broader coverage, as it encompasses the experimental settings of the current GSM8K data settings (i.e., 1-shot). For the task-aware dataset, we use NaturalQuestions (Liu et al., 2023). The statistics of dataset is given in Table 5, and their detailed description is provided in Appendix D.1.
Baselines. We compare the performance of Prompt-SAW against following models as baselines: (i) Selective-Context (Li, 2023), (ii) LLMLingua (Jiang et al., 2023a), (iii) LongLLMlingua (Jiang et al., 2023b), and (iii) GPT4 (Achiam et al., 2023). Details about the baselines are provided in Appendix D.2.
Evaluation Metrics. For GSM8k-aug, we use Exact Match (EM) as the evaluation metric. This metric is also employed by (Cobbe et al., 2021) and (Jiang et al., 2023a). For the evaluation of NaturalQuestions data, we used Span Accuracy (Span-Acc) as a metric. This is similar to previous work by (Liu et al., 2023) and (Jiang et al., 2023b). Apart from this, we also use fluency (FL) (Meng et al., 2022) to measure the readability and grammatical coherence of the compressed prompt. The details about these metrics are given in Appendix D.3.
(GSM8k-aug)
(1-shot) (2-shot) (4-shot)
Method EM Tokens $1/\eta$ EM Tokens $1/\eta$ EM Tokens $1/\eta$
Original No.1 (1-shot) 71.04 422 1.00
No.2 (1-shot) 74.98 311 1.00 76.81 733 1.00
No.3 (1-shot) 73.92 249 1.00
No.4 (1-shot) 73.39 242 1.00 79.53 491 1.00 78.92 1224 1.00
Selective-Context No.1 (1-shot) 53.28 317 1.33
No.2 (1-shot) 54.74 231 1.35 57.34 550 1.33
No.3 (1-shot) 58.40 190 1.32
No.4 (1-shot) 55.78 175 1.39 52.17 349 1.41 58.13 881 1.39
LLMLLingua No.1 (1-shot) 68.43(0.5%) 321 1.32
No.2 (1-shot) 60.02 234 1.33 65.71 498 1.47
No.3 (1-shot) 66.53 202 1.23
No.4 (1-shot) 62.38(4.7%) 172 1.41 59.29 358 1.37 63.41(13.7%) 906 1.35
GPT4-Generation No.1 (1-shot) 62.86 283 1.49
No.2 (1-shot) 65.71(9.5%) 221 1.41 68.57(4.3%) 469 1.56
No.3 (1-shot) 66.57(1.1%) 199 1.25
No.4 (1-shot) 60.18 159 1.52 67.00(8.2%) 354 1.39 62.68 956 1.28
Prompt-SAW(Ours) No.1 (1-shot) 68.78 292 1.45
No.2 (1-shot) 71.98 212 1.47 71.53 469 1.56
No.3 (1-shot) 67.27 160 1.56
No.4 (1-shot) 65.32 165 1.47 72.47 329 1.49 72.12 820 1.49
Table 1: Experimental results on GSM8k-aug. We report the number of tokens in original and compressed prompts along with EM and compression ratio $(1/\eta)$. We bold-face the overall best-performing scores with existing state-of-art underlined.
Experimental Setup. In this work, we select GPT-3.5-Turbo-0613 as the target LLM, which is accessible via OpenAI API.  [^1] https://openai.com/blog/openai-api Following the setting of LLMLingua (Jiang et al., 2023a), we employ greedy decoding with the temperature set to 0. For graph construction, we use open-source language model  Vicunna-7B (Chiang et al., 2023). We use OpenAI embedding API  [^2] https://openai.com/ as embedding encoder $(E)$. The value for $\eta^{*}$ is set to {0.1, 0.2, 0.3, 0.4, 0.5} for both  GSM8k-aug and NaturalQuestions. Value of $\gamma$ in Algorithm 2 is 0.001. All the results reported in this paper are averaged over five runs. All experiments were performed using PyTorch 2.1.0 with Nvidia RTX 4090 24GB GPU.
5.2 Experimental Results
Results for Task-agnostic Settings. We first study the performance of Prompt-SAW for task-agnostic settings, whose results on GSM8k-aug are shown in Table 1. Unlike existing research that reports their performance for the 1-shot setting, we report these results for 1-shot, 2-shot and 4-shot settings, where $i$-shot indicates $i$ numbers of prompts have been employed by Prompt-SAW.
Comparing these results against the baseline models, we can observe that Prompt-SAW outperforms the previous state-of-the-art by a significant margin. For instance, compared to the best performing baselines, Prompt-SAW improves the EM score by up to 9.5%, 8.2%, and 13.7% under 1-shot, 2-shot and 4-shot settings, respectively. Correspondingly reduction in the prompt size is 31.8%, 33.0%, and 33.0%. We attribute such drastic performance improvement to the following factors: (1) Prompt-SAW retains the logical integrity of the COT prompts by sub-dividing the original prompts into smaller comprehensive information elements; (2)Prompt-SAW benefits from the workflow that allows selecting and omitting individual information elements for prompt compression without destroying the overall information structure of the compressed prompt. These help Prompt-SAW to ensure the utility of the compressed prompt for the end task.
Results for Task-aware Settings. Table 2 reports the performance of Prompt-SAW under task-aware settings on NaturalQuestions. These results show that, Prompt-SAW improves the Span Accuracy by 14.30%, 10.62% and 6.52%, respectively, for different values of the target compression rates, i.e., $\eta^{*}=\{0.5,0.3,0.1\}$, against the best-performing baseline (GPT-4). Correspondingly, the reduction in the prompt size is 56.7%, 74.0%, and 93.7%.
Correlating the results for both settings, we observe that compared to the task-agnostic scenarios, Prompt-SAW yields much stable performance for the task-aware settings. This is due to the fact that it is more difficult to dig out the latent correlation between information elements within the prompt’s internal structure rather than explicit task-aware correlation extraction. From Table 2 and Figure 3 we can also find that the performance of Prompt-SAW drops when the value for the target compression rate $(\eta^{*})$ decreases from 0.5 to 0.1. A probable justification that the actual compression ratio of Prompt-SAW is significantly higher than the target compression ratio when the target compression ratio is 10 (i.e., $\eta^{*}=0.1$). This is owing to the fact that Prompt-SAW only retains the information elements in $\mathcal{G}$ as the key/basic information units for the compression process. It will delete some entities and relations that may be highly similar to the problem but their overall structure is too long, leading to relatively poor performance for higher compression ratio.
Method NaturalQuestions
$\eta^{*}$ = 0.5 $\eta^{*}$ = 0.3 $\eta^{*}$ = 0.1
Span-Acc Tokens 1/$\eta$ Span-Acc Tokens 1/$\eta$ Span-Acc Tokens 1/$\eta$
Original 92.18 524 1.00 92.18 524 1.00 92.18 524 1.00
Selective-Context 49.23 283 1.85 45.71 173 3.03 31.12 68 7.69
LongLLMlingua 59.65 270 1.94 52.02 161 3.25 47.14 57 9.21
GPT4-Generation 72.55(14.30%) 252 2.08 66.19(10.62%) 121 4.35 50.76(6.52%) 84 6.25
Prompt-SAW(Ours) 82.93 227 2.31 73.22 136 3.86 54.07 33 16.08
Table 2: Experimenal results on NaturalQuestions.
Target LLMs GSM8K-AUG Natural Questions
1shot 2shot 4shot
GPT3.5 67.83 72.00 72.12 73.22
GPT4 91.48 93.17 95.92 74.36
Table 3: The performance comparison of Prompt-SAW using different target LLMs.
5.3 Further Analysis
Different Target LLMs. We also analyze the performance of Prompt-SAW using different target LLMs, with results reported in Table 3. We can find that GPT4 performs better than GPT3.5 under the task-agnostic setting, while they perform similarly under the task-aware setting. We attribute this result to GPT4’s stronger inference capabilities, so it can benefit more and learn how to perform mathematical reasoning from compressed prompts in a task-agnostic setting. However, when we consider the task-aware settings, the performance of GPT4 and GPT3.5 is similar. That is because the LLMs need to extract key information from the prompt to answer the question. If we compress the important information, LLM will make mistakes no matter how strong the reasoning ability it has.
Refer to caption Refer to caption
Figure 3: Different target compress rate between LLMLingua, GPT4 and Prompt-SAW.
Interpretability of Compressed Prompts. As explained in the introduction (also highlighted in Figure 1), a key limitation of existing prompt compression approaches is the limited interpretability of the compressed prompt. In order to validate the results of Prompt-SAW in terms of human readability and/or interpretability, we report some example prompts along with prompt compressed using Prompt-SAW and LLMLingua (Jiang et al., 2023a) in Appendix E.1, for a quick comparison. These examples clearly indicate that the prompt compressed by Prompt-SAW exhibit better readability and/or interpretability compared to compressed using LLMLingua. This loss of interpretability may significantly impact the end-utility of the compressed prompt.
We also conducted a quantitative comparison to support our claims. Specifically, we assess the fluency of the compressed prompts through the computation of a weighted mean of bi-gram and tri-gram entropies (Meng et al., 2022). Corresponding results in Table 4 show that Prompt-SAW yields relatively higher fluency scores than the baseline models. A lower score for baseline models, e.g., LLMLingua, may be attributable to loss of intrinsic semantic relationship between the tokens for the compressed prompt.
Selective-Context LLMLingua GPT4 Prompt-SAW
FL 5.61 5.74 6.07 6.30
6 Conclusion
In this work, we propose Prompt-SAW that leverages graph structures to infer key information in the prompt in order to come up with a compressed prompt. Experimental evaluation shows that Prompt-SAW outperforms the existing research on prompt compression by a significant margin. Moreover, Prompt-SAW also addressed a key limitation of existing prompt compression approaches, i.e., the compressed prompts are human-friendly and easy to understand. In the future, we also plan to extend this work to other languages by leveraging techniques for mapping embedding vectors (Ali et al., 2023a; b).
Appendix A Background
A.1 Knowledge Graph
A.2 Task-aware Prompts
A.3 Task-agnostic Prompts
A.4 Chain-of-thought Prompt
A.5 Prompt with external documentation
A.6 Token-level Prompt compression
LLMLingua is a token-level prompt compression method that performs compression based on perplexity. It includes a budget controller to calculate the compression ratio of demonstrations and the further compression ratio of each demonstration based on given parameters. LLMLingua uses the LLAMA-2 (Touvron et al., 2023) to calculate the perplexity of each token. Finally, LLMLingua compresses the prompt based on the perplexity and compression ratio.
Appendix B Prompts
B.1 Prompts for Graph Construction
Prompts for Graph Construction: Example:
Input:
Output:
B.2 Instructions used for GPT-4 response Generation
The instructions we used in the GPT-4 Generation are shown below:
Appendix C Additional Implementation details
Algorithm 3 Compress Task-agnostic Prompt

### v2

@@ Abstract
Large Language Models (LLMs) have shown exceptional abilities for multiple different natural language processing tasks. While prompting is a crucial tool for LLM inference, we observe that there is a significant cost associated with exceedingly lengthy prompts. Existing attempts to compress lengthy prompts lead to sub-standard results in terms of readability/interpretability of the compressed prompt, with a detrimental impact on prompt utility. To address this, we propose Prompt-SAW: Prompt compresSion via Relation AWare graphs, an effective strategy for prompt compression over task-agnostic and task-aware prompts. Prompt-SAW uses the prompt’s textual information to build a graph, later extracts key information elements in the graph to come up with the compressed prompt. We also propose GSM8K-aug, i.e., an extended version of the existing GSM8K benchmark for task-agnostic prompts in order to provide a comprehensive evaluation platform. Experimental evaluation using benchmark datasets shows that prompts compressed by Prompt-SAW are not only better in terms of readability, but they also outperform the best-performing baseline models by up to 10.1% and 77.1% respectively for task-agnostic and task-aware settings while compressing the original prompt text by 34.9% and 56.7%.
@@ 1 Introduction
LLMs have attracted considerable attention for their superior performance across a wide range of applications. For this, instructions (aka. prompts) play a crucial role in extending the capabilities of LLMs for multiple different tasks. The prompts provide the provision to guide the model to elucidate desired model behavior without perturbing the model parameters. This is also highlighted in recent studies that show well-designed prompts and the integration of external knowledge are significant to enhance the effectiveness of LLMs’ (Sahoo et al., 2024). Different LLMs-related techniques directly benefiting from prompts include but are not limited to: In-Context Learning (Dong et al., 2022), Chain-of-Thought (Wei et al., 2022), Retrieval Augmented Generation (Lewis et al., 2020), and Agents (Park et al., 2023) etc. Generally, prompts may be sub-divided into two types: task-aware and task-agnostic prompts, a quick overview is given in Appendix A.2 and Appendix A.3 respectively.
At the same time, the abilities of LLMs are significantly compromised/constrained by increasingly lengthy prompts, even comprising thousands of tokens. Lengthy prompts not only obscure requisite information but also increase computational costs and incur inference latency. To tackle this challenge, prompt compression techniques, e.g., Li (2023), have garnered significant interest. These approaches are based on the fact that natural language is inherently redundant (Shannon, 1951). Thus, it is possible to substantially compress the length of original textual prompts by preserving requisite information in small segments.
@@
While existing approaches could enhance the ability to deal with lengthy prompts for LLMs, they lack grammatical coherence, i.e., existing approaches neglect the syntactic and semantic structure of the compressed prompt. This is because contemporary prompt compression methods primarily focus on quantifying token-level information, neglecting the overall grammatical structure of the compressed prompt. Such ignorance not only increases the risk of semantic loss within the compressed prompt but also hampers its readability for human readers. An example in this regard is shown in Figure 1, where the original prompt text: ”Two women have won the prize: Curie and Maria Goeppert-Mayer” is compressed to: ”won twoes forg01 theate women prize:ertMayer” by LongLLMlingua (Jiang et al., 2023b).
Figure 1: An example of compressed prompt compressed using previous token-level based method LongLLMlingua (Jiang et al., 2023b)
@@
We conducted extensive experimental analysis of Prompt-SAW under both task-agnostic and task-aware settings against existing best-performing models as baselines. For evaluation, we used: (i)  GSM8K-aug, i.e., an extended experimental setting proposed by us for GSM8K (Cobbe et al., 2021), (ii) NaturalQuestions (Liu et al., 2023), and (iii) ShareGPT [^1] https://sharegpt.com/. Experimental results show that Prompt-SAW significantly outperforms other baseline models. We summarize the key contributions of this work as follows:
@@
  • As current benchmarks for task-agnostic prompts lack comprehensive evaluation, we propose GSM8K-aug, an extended version of the existing GSM8K benchmark for an intensive evaluation of Prompt-SAW.
  • We demonstrate the effectiveness of Prompt-SAW by comprehensive experiments, showing Prompt-SAW attained state-of-the-art performance outperforming baseline models by up to 10.1% and 77.1% respectively for task-agnostic and task-aware settings while compressing the original prompt text by 34.9% and 56.7%.
@@ 2 Related Work
Soft prompts were introduced by Lester et al. (2021). A soft prompt integrates additional trainable parameters at the model’s input stage. Wingate et al. (2022) emphasized that soft prompt compression effectively retains crucial abstract information with a reduced parameter count. Xu et al. (2023) emphasized that carefully crafted prompts are helpful in augmenting the end-performance of compressed LLMs, also the compressed LLMs are helpful in the prompt learning phase.
Compared to soft prompt compression, discrete prompt compression seeks to optimize the effectiveness of prompts via token-level search strategies. Jung and Kim (2023) employed policy networks to eliminate unnecessary tokens for prompt compression. Li (2023) utilized self-information metrics to identify and remove superfluous information in prompts. Capitalizing on these advancements, Jiang et al. (2023a) and Jiang et al. (2023b) formulated algorithms for dynamically adjusting compression rates across different prompt sections, giving precedence to tokens with higher perplexity.
@@ 3 Preliminaries
We use $P$ and $C$ to represent the original and compressed prompt respectively. Likewise, we use $N$ and $\widetilde{N}$ to represent the length of the original and compressed prompt. We use $\eta=\widetilde{N}/N$ to represent the compression rate and $1/\eta$ as the compression ratio. $\eta^{*}$ is used to represent the target compression rate. The graph is represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $e_{i}$, $r_{i}$ and $e_{i}^{{}^{\prime}}$ represent the subject entity, relation and object entity in the graph respectively; =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ and =$\mathcal{R}=\{r_{1},r_{2},\cdots,r_{n}\}$ denote the set of entities and relations in $\mathcal{G}$. $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is used to represent small-scale information elements in $\mathcal{G}$, equivalent to graph triplet. $M$ represents auxiliary models used for graph construction. $E$ is used to represent the encoder network. $\delta$ represents the similarity threshold used for sub-graph construction.
@@ 4.1 Motivation of Prompt-SAW
Formally, given a prompt text $P$, we claim it encompasses a set of entities $\mathcal{E}$, i.e., names of persons, locations, organizations, miscellaneous elements, etc., (Ali et al., 2020). These entities serve as the key elements of the prompt structure. In addition to the entities, we can also infer some relations $\mathcal{R}$ in $P$ that may be used to describe the connections between the entities. Prompt-SAW re-organizes these key elements of the prompt (i.e., entities and their relations) in a graph structure, represented by $\mathcal{G}=\{(e_{i},r_{i},e_{i}^{{}^{\prime}})\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$. We use $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ to represent the $i$-th information element of $\mathcal{G}$, i.e., a fact stating that $e_{i}$ has $r_{i}$-$th$ relation with $e_{i}^{{}^{\prime}}$.
@@ 4.2 Workflow of Prompt-SAW
The workflow of Prompt-SAW consists of two parts. First, it uses the information in prompt $P$ to construct a graph $\mathcal{G}$. Then, based on the specific scenario, we proceed as follows:
@@
Further details about the model components of Prompt-SAW are provided in the following subsections.
4.2.1 Graph Construction.
For graph construction from the text data, we primarily rely traditional knowledge extraction approaches, i.e., OpenIE (Kolluru et al., 2020), to construct a graph $G$, as follows
$\mathcal{G}={IE}({P}),$ (1)
where $P$ is the prompt text, and $IE$ is an information extraction module that takes $P$ as input and return graph $G$ as output. For the cases not addressed by the above equation, we use an in-context learning prompt as our auxiliary method that prompts the language model to construct the graph from the original prompt text as follows:
$\mathcal{G}={M}(P_{\text{template}}(\text{P})),$ (2)
where $(P)$ is the prompt text and $P_{\text{template}}$ is the prompt template (explained in Appendix B.1) used to guide the LLM $(M)$ to extract the graph $\mathcal{G}$. Note, for Equation 2, we typically prefer a small-scale open-source LLM in order to avoid higher computational costs incurred by large models.
4.2.2 Task-aware Prompts
$Emb_{p^{\mathrm{que}}}=E(p^{\mathrm{que}})$ (3)
$Index_{\text{ranked}}=\text{Rank}(Sim_{\mathcal{G}})$ (5)
$C=e_{1}\oplus r_{1}\oplus e_{1}^{\prime};\cdots;e_{n}\oplus r_{n}\oplus e_{n}^{\prime}$ (6)
where $\oplus$ is the concatenation operator used to combine the entities and relations within the information elements $(g_{i})$ in the extracted subgraph $\mathcal{G}_{\text{subset}}$, and (;)({;})( ; ) is the delimiter used to separate different information elements in $\mathcal{G}_{\text{subset}}$.
@@ 4.2.2 Task-aware Prompts
4.2.3 Task-agnostic Prompts
Compress Function. The workflow of the $Compress()$ is shown in Algorithm 3 and explained as follows: (i) start with an empty graph $(\mathcal{G}^{\prime})$ (line-1); (ii) iterate the information elements in the graph $(g_{i}\in\mathcal{G})$ to compute the similarity score of $g_{i}$ with the elements in $\mathcal{G}^{\prime}$ to look for maximal similarity, i.e., $sim_{max}$ (lines 2-3); (iii) compare $sim_{max}$ against the compression threshold $\delta$ to insert $g_{i}$ in $(\mathcal{G}^{\prime})$ (lines 4-5); (iv) finally, return $(\mathcal{G}^{\prime})$ as the final subset of the graph. The end-goal of Algorithm 2 is to select highly dis-similar information elements by neglecting cases with $sim_{max}>\delta$. For such cases, we assume that the corresponding information element, i.e., $g_{i}=(e_{i},r_{i},e_{i}^{{}^{\prime}})$ is redundant because there is already an element in $\mathcal{G}^{\prime}$ that is very similar to $g_{i}$.
@@ 4.2.3 Task-agnostic Prompts
5 GSM8K-AUG
As a benchmark dataset, GSM8K (Cobbe et al., 2021) encompasses high-quality, linguistically diverse grade math word problems. However, we find that the original dataset poses some limitations for evaluating the prompt compression methods. Specifically, it only allows compressing prompts under one fix setting, i.e., 8-shot. This is inadequate for rigorously evaluating the abilities of the prompt compression systems. For instance, it makes it harder to analyze and answer the questions: (i) Whether prompt compression methods destroy connections between individual shots? (ii) Also, what impact will these connections have on the end-performance for in-context-learning tasks?
To address these limitations, we propose GSM8K-aug, an extended and more comprehensive experimental setting for original GSM8K data. GSM8K-aug extends the original data set to $i$-shot setting ($i\in\{1,2,4,8\}$), with $i$-shot meaning $i$ example demonstrations in the prompt. Note, GSM8K-aug has a broader coverage, as it encompasses the experimental settings of the current GSM8K data settings (i.e., 8-shot). We argue GSM8K-aug helps in overcoming the limitations mentioned above, as it provides us with the provision to find correlations between different shots. For instance, it can help us to quickly answer the above questions by analyzing the models’ performance by compressing two prompts at the same time, i.e., 2-shot settings compared against compressing them independently, i.e., 1-shot settings.
Algorithm 3 Compress Prompt